# Enhancing LLM-based Test Generation for Hard-to-Cover Branches via Program Analysis

### v1

@@ Abstract
Automatic test generation plays a critical role in software quality assurance. While the recent advances in Search-Based Software Testing (SBST) and Large Language Models (LLMs) have shown promise in generating useful tests, these techniques still struggle to cover certain branches. Reaching these hard-to-cover branches usually requires constructing complex objects and resolving intricate inter-procedural dependencies in branch conditions, which poses significant challenges for existing techniques. In this work, we propose TELPA, a novel technique aimed at addressing these challenges. Its key insight lies in extracting real usage scenarios of the target method under test to learn how to construct complex objects and extracting methods entailing inter-procedural dependencies with hard-to-cover branches to learn the semantics of branch constraints. To enhance efficiency and effectiveness, TELPA identifies a set of ineffective tests as counter-examples for LLMs and employs a feedback-based process to iteratively refine these counter-examples. Then, TELPA integrates program analysis results and counter-examples into the prompt, guiding LLMs to gain deeper understandings of the semantics of the target method and generate diverse tests that can reach the hard-to-cover branches. Our experimental results on 27 open-source Python projects demonstrate that TELPA significantly outperforms the state-of-the-art SBST and LLM-based techniques, achieving an average improvement of 31.39% and 22.22% in terms of branch coverage.
I Introduction
@@
Automatic test generation holds significant importance in software quality assurance, which aims to efficiently detect software bugs by automatically covering various behaviors of the software under test. Recently, many test generation techniques have been proposed, among which Search-Based Software Testing (SBST) [1, 2, 3] is one of the most widely-studied categories due to its effectiveness. Typically, SBST techniques leverage heuristic search algorithms to explore the vast test space by mutating already-generated tests with the guidance of maximizing code coverage. While receiving extensive attention, they are still unable to reach specific parts of code. For instance, they often fail to generate tests to cover the branches that can be triggered only by some scenario-specific values (e.g., specific file names, version strings) due to the lack of the ability of understanding code semantics [4].
@@
The recent advancement of Large Language Models (LLMs) provides new opportunities for tackling these problems, and several LLM-based techniques have been proposed to generate effective test cases [4, 5]. These techniques utilize the code comprehension ability of LLMs by incorporating the source code of the target method to be tested and some contextual information (such as other methods around the target method) into the prompt. While the limitation of SBST can be relieved to some degree with the help of LLMs, the automatically generated tests still tend to be generic and only able to cover the branches without complicated constraints [6]. That is, many hard-to-cover branches are still left for manual test design based on experts’ domain knowledge, leading to significant costs. An automated approach which can further enhance the coverage of these hard-to-cover branches can largely reduce the testing effort.
@@
In practice, two major challenges stand out when creating such approaches. First, some branch constraints involve objects with complex construction processes. Specifically, the construction of some objects relies on other (complex) types of objects, ultimately necessitating a specific sequence of constructor invocations. As indicated by a previous study [5], the tests generated by LLMs exhibit a relatively low compilation success rate (about 39%), highlighting a deficiency in the ability of LLMs in this regard. Hence, state-of-the-art LLM-based techniques (e.g., CODAMOSA [4]) fail to produce valid objects for such scenarios, let alone the objects with specific attribute values required by branch constraints. Second, some branch constraints entail intricate inter-procedural dependencies. That is, the outcome of certain conditions within the constraints are determined by the execution results of a series of method invocations. In such scenarios, merely providing source code of the target method or some coarse-grained contextual information (the common practice of existing LLM-based techniques) is insufficient to enable LLMs to comprehend the semantics conveyed by a sequence of invoked methods, thereby struggling to generate effective tests.
@@
To tackle these challenges, we propose a novel LLM-based test generation technique (called TELPA – TEst generation via Llm and Program Analysis), which focuses on reaching hard-to-cover branches via program-analysis-enhanced prompting. Given that complex objects in branch constraints are often passed to the target method as parameters, TELPA addresses the first challenge by gathering invocations of the target method within the software module. These existing invocations might contain the whole process of constructing these objects. Specifically, TELPA conducts backward method-invocation analysis, starting from each invocation of the target method, to extract all sequences of method invocations with the target method as the endpoint. Then, TELPA prompts LLMs to generate tests by using different paths to enter the target method. By exploring different entry paths, TELPA not only facilitates the construction of valid objects but also captures various usage scenarios in practice (that may produce specific attribute values required by branch constraints), thus enhancing test diversity and improving coverage.
@@
To address the second challenge arising from complicated inter-procedural dependencies in branch constraints, TELPA performs forward method-invocation analysis. This analysis begins from each variable and invoked method in the constraints for a target branch, aiming to extract all associated methods. By incorporating the source code of these methods in their invocation order as the prompt, TELPA can elicit LLMs to gain a deeper understanding of the semantics pertaining to the target branch, thereby generating effective tests. This provides precise contextual information concerning the target branch and it is more effective than directly supplying the source code of all methods within the software module to LLMs, as the latter could introduce excessive noise irrelevant to the target branch and confuse LLMs during test generation.
@@
Due to the non-negligible cost of utilizing LLMs, we define the usage scenario of TELPA as its activation when the existing (lightweight) test generation tool (e.g., Pynguin [1]) fails to increase test coverage within a certain timeframe. In other words, TELPA is only employed for hard-to-cover branches to ensure cost-effectiveness. To enhance efficiency further, TELPA samples a diverse set of already-generated tests as counter-examples. By incorporating these counter-examples into the prompt, TELPA can instruct LLMs to generate tests that diverge from them, as these counter-examples have been recognized as ineffective for these hard-to-cover branches. In particular, the overall test generation process with TELPA is structured as a feedback-based process. This design allows for updating counter-examples based on the most recent coverage results, thereby enhancing test generation effectiveness.
@@
We conducted an extensive evaluation on 27 open-source Python projects, which have been widely used in previous studies [4, 1]. Our results show that TELPA significantly outperforms both Pynguin (a state-of-the-art SBST technique) and CODAMOSA (a state-of-the-art LLM-based technique). On average, TELPA achieves 31.39% and 22.22% higher branch coverage than them respectively, given the same testing time budget. Our ablation study also confirms the contribution of each main component in TELPA, including backward method-invocation analysis for relieving the challenge arising from complex object construction, forward method-invocation analysis for relieving the challenge arising from intricate inter-procedural dependencies, counter-examples sampling and coverage-based feedback for improving the overall testing efficiency. Our investigation into various configurations of TELPA confirms that our designed usage scenario for TELPA strikes a balance between effectiveness and efficiency and TELPA consistently achieves higher coverage than other settings.
@@
We introduce TELPA, a novel LLM-based test generation technique specifically designed to enhance the coverage of hard-to-cover branches through program-analysis-enhanced prompting.
TELPA can well address the challenges arising from complex object construction and intricate inter-procedural dependencies by incorporating both backward and forward method-invocation analysis within TELPA.
@@
We present an extensive evaluation of TELPA across 27 open-source projects, demonstrating its effectiveness in comparison to a state-of-the-art SBST technique and a state-of-the-art LLM-based one.
II Motivating Example
In this section, we demonstrate the limitations of the state-of-the-art techniques Pynguin [1] and CODAMOSA [4] with two real examples to motivate our study.
Figure 1 shows the simplified code snippet of the target method set_definition from the typesystem project111https://github.com/encode/typesystem. With a time budget of 20 minutes, both Pynguin and CODAMOSA fail to reach line 6, namely the condition for the if statement in line 5 never returns true. In fact, to make the condition true, the generated test needs to first construct an Array object (line 3), and then assign a proper value to its items attribute. More specifically, items should be either a tuple or a list of Field objects (line 4), according to the type hints from the constructor of Array (line 13). Constructing such a sophisticated Array object is certainly challenging for automatic test generation techniques, Pynguin and CODAMOSA are no exception.
Another type of hard-to-cover branches are caused by complicated inter-procedural dependencies, as illustrated in Figure 2. In this case, our target method is is_public_family from the apimd project222https://github.com/KmolYuan/apimd. Both Pynguin and CODAMOSA are unable to guide the condition in line 3 to return true. The challenge here is that the outcome of the if condition depends on the method is_magic, which is defined elsewhere in the project and might further invoke other methods. Those complicated inter-procedural dependencies are often hard to understand and analyze for test generation techniques.
@@ II Motivating Example
First, during the test generation process, existing techniques [1, 4] normally try to construct minimal or over-simplified objects, without taking into account the real usage scenarios involving specific attributes. Second, for the inter-procedural dependencies, state-of-the-art techniques [1, 4] are not able to recursively analyze how dependent methods interact with variables from the target method, thus failing to interpret the concrete branch constraints. Third, existing LLM-based techniques [4, 5] typically provide the whole module as the context for LLMs to generate tests. In reality, many parts of code is irrelevant to the target method, which can only introduce noise to LLMs. Moreover, a prompt can easily exceed the input length limit if the information of the whole module is incorporated, resulting in undesired prompt truncation and critical information loss.
@@
Based on these potential reasons behind the unsatisfactory performance of existing techniques, we propose TELPA, an LLM-based test generation technique leveraging program-analysis-enhanced prompting. TELPA will 1) conduct backward method-invocation analysis to learn real usage scenarios of objects and overcome the difficulty of constructing complex objects, 2) conduct forward method-invocation analysis to understand the behavior of complicated inter-procedural dependencies, and 3) only supply relevant information to LLMs for test generation, avoiding information interference.
III Approach
@@
The overall workflow of TELPA is illustrated in Figure 3. First, TELPA leverages existing test generation tools to generate a set of tests, which can reach those easy-to-cover branches. For ease of presentation, we call this step preceding test generation. Then, for the methods containing the branches that are still not covered, TELPA performs two types of program analyses, including backward method-invocation analysis to retrieve the method invocation sequences for complex object construction and forward method-invocation analysis to retrieve all associated methods for inter-procedural dependencies. To improve the testing efficiency, TELPA samples existing generated tests as counter-examples to guide LLMs to generate different tests, as they are shown ineffective to cover specific branches. These counter-examples, together with the program analysis results, are incorporated into the prompt for LLMs to generate new tests. The new tests will be executed and added into existing tests for future iterations.
III-A Preceding Test Generation with Existing Tools
@@
As our goal is to reach those hard-to-cover branches, TELPA outsources the initial testing task (for those easy-to-cover branches) to the existing test generation tools. This helps to improve the cost-effectiveness of the overall testing process due to the cost of utilizing LLMs. Specifically, for a project under test, TELPA runs selected state-of-the-art test generation tools until they are unable to cover new branches within a pre-defined timeframe. When executing these generated tests, their corresponding coverage information is also recorded. It is worth noting that TELPA is in theory language-independent and not restricted to specific tools for generating initial tests. In this section, we mainly describe the idea behind TELPA, and the implementation details will be presented in Section IV-A1.
III-B Method-Invocation Analyses
@@
TELPA first collects the methods containing branches not covered by the preceding testing process, and then performs backward and forward method-invocation analyses on these methods to gather information needed for handling complex object construction and inter-procedural dependencies. This process is done in ascending order of coverage achieved for the methods, as we assume that methods with the lower coverage are more difficult to tackle and there is larger room for coverage improvement. Unlike existing LLM-based techniques (e.g., CODAMOSA [4]) that try to incorporate as much source code from the module as possible into prompts, we will only feed LLMs with the methods retrieved from method-invocation analyses in order to reduce noise introduced by irrelevant code and alleviate computational burden of LLMs.
III-B1 Backward Analysis

### v2

# Advancing Code Coverage: Incorporating Program Analysis with Large Language Models

@@ Abstract.
Automatic test generation plays a critical role in software quality assurance. While the recent advances in Search-Based Software Testing (SBST) and Large Language Models (LLMs) have shown promise in generating useful tests, these techniques still struggle to cover certain branches. Reaching these hard-to-cover branches usually requires constructing complex objects and resolving intricate inter-procedural dependencies in branch conditions, which poses significant challenges for existing techniques. In this work, we propose TELPA, a novel technique aimed at addressing these challenges. Its key insight lies in extracting real usage scenarios of the target method under test to learn how to construct complex objects and extracting methods entailing inter-procedural dependencies with hard-to-cover branches to learn the semantics of branch constraints. To enhance efficiency and effectiveness, TELPA identifies a set of ineffective tests as counter-examples for LLMs and employs a feedback-based process to iteratively refine these counter-examples. Then, TELPA integrates program analysis results and counter-examples into the prompt, guiding LLMs to gain deeper understandings of the semantics of the target method and generate diverse tests that can reach the hard-to-cover branches. Our experimental results on 27 open-source Python projects demonstrate that TELPA significantly outperforms the state-of-the-art SBST and LLM-enhanced techniques, achieving an average improvement of 34.10% and 25.93% in terms of branch coverage.
1. Introduction
@@
Automatic test generation holds significant importance in software quality assurance, which aims to efficiently detect software bugs by automatically covering various behaviors of the software under test. Recently, many test generation techniques have been proposed, among which Search-Based Software Testing (SBST) (Lukasczyk and Fraser, 2022; Pacheco and Ernst, 2007; Fraser and Arcuri, 2011a) is one of the most widely-studied categories due to its effectiveness. Typically, SBST techniques leverage heuristic search algorithms to explore the vast test space by mutating already-generated tests with the guidance of maximizing code coverage. While receiving extensive attention, they are still unable to reach specific parts of code. For instance, they often fail to generate tests to cover the branches that can be triggered only by some scenario-specific values that may require deep understanding of the code semantics.
@@
The recent advancement of Large Language Models (LLMs) provides new opportunities for tackling these problems, and several LLM-based techniques have been proposed to generate effective test cases (Lemieux et al., 2023; Yuan et al., 2024). These techniques utilize the code comprehension ability of LLMs by incorporating the source code of the target method to be tested and some contextual information (such as other methods around the target method) into the prompt. While the limitation of SBST can be relieved to some degree with the help of LLMs, the automatically generated tests still tend to be generic and only able to cover the branches without complicated constraints (Ryan et al., 2024). That is, many branches are still left for manual test design based on experts’ domain knowledge, leading to significant costs. An automated approach that can further improve the coverage of these branches can largely reduce the testing effort.
@@
In practice, two major challenges stand out when creating such approaches. First, some branch constraints involve objects with complex construction processes. Specifically, the construction of some objects relies on other (complex) types of objects, ultimately necessitating a specific sequence of constructor invocations. For example, the branch condition at Line 5 in Figure 1 requires a field object with a specific type (e.g., Array) and a specific attribute value (e.g., items should be a (tuple, list) with the type typing.Union[Field, typing.Sequence[Field]]). Constructing such a field object is challenging because it requires understanding the dependencies between multiple objects: (1) constructing valid items (which must be either a single Field object or a sequence of Field objects), and (2) creating a valid field object, ensuring that the items attribute has the correct type. This process involves multiple steps and dependencies that are difficult for existing techniques to handle without a deep understanding of the code semantics. As indicated by a previous study (Yuan et al., 2024), the tests generated by LLMs exhibit a relatively low compilation success rate (about 39%), which directly impacts their effectiveness in achieving code coverage. This highlights a deficiency in the ability of LLMs to generate effective tests in such scenarios. Indeed, state-of-the-art LLM-based techniques (e.g., CODAMOSA (Lemieux et al., 2023)) fail to produce valid objects for such scenarios, let alone the objects with specific attribute values required by branch constraints. Second, some branch constraints entail intricate inter-procedural dependencies. That is, the outcome of certain conditions within the constraints are determined by the execution results of a series of method invocations. For instance, the branch condition at Line 3 in Figure 2 depends on the result of the method is_magic. This method is defined elsewhere in the project and may invoke other methods, creating a chain of inter-procedural dependencies. To cover this branch, the test must correctly simulate the action that is_magic performs and provide the test data that satisfy the condition. This requires a deep understanding of the code semantics to ensure that the generated test accurately reflects the dependencies and produces the correct result for the branch condition. In such scenarios, merely providing source code of the target method or some coarse-grained contextual information (the common practice of existing LLM-based techniques) is insufficient to enable LLMs to comprehend the semantics conveyed by a sequence of invoked methods, thereby struggling to generate effective tests. Therefore, we define “hard-to-cover branches” as those that meet either of the following criteria: (1) branches requiring values derived from complex object construction. For example, if a branch condition checks whether an object’s attribute matches specific conditions (such as type, format, and other non-constant conditions), the test must first construct the object with the appropriate attribute to cover the branch; or (2) branches with complex inter-procedural dependencies. For example, if a branch condition depends on the result of a method invocation, which itself relies on other method invocations, the test must accurately simulate the sequence of method calls and their interactions to cover the branch.
@@
To tackle these challenges in covering hard-to-cover branches, we propose a novel LLM-based test generation technique (called TELPA – TEst generation via Llm and Program Analysis), which focuses on reaching hard-to-cover branches via program-analysis-enhanced prompting. Given that complex objects in branch constraints are often passed to the target method as parameters, TELPA addresses the first challenge by gathering invocations of the target method within the software module. These existing invocations might contain the whole process of constructing these objects. Specifically, TELPA conducts object construction analysis, which traces the sequence of method calls leading to the target method, focusing on understanding how objects are constructed and passed to the target method. This analysis starts from each invocation of the target method, to extract all sequences of method invocations with the target method as the endpoint. Then, TELPA prompts LLMs to generate tests by using different paths to enter the target method. By exploring different entry paths, TELPA not only facilitates the construction of valid objects but also captures various usage scenarios in practice (that may produce specific attribute values required by branch constraints), thus enhancing test diversity and improving coverage.
@@
To address the second challenge arising from complicated inter-procedural dependencies in branch constraints, TELPA performs branch dependency analysis, which determines the methods invoked within the branch conditions of the target method, focusing on understanding the inter-procedural dependencies that influence branch outcomes. This analysis begins from each variable and invoked method in the constraints for a target branch, aiming to extract all associated methods. By incorporating the source code of these methods in their invocation order as the prompt, TELPA can elicit LLMs to gain a deeper understanding of the semantics pertaining to the target branch, thereby generating effective tests. This provides precise contextual information concerning the target branch and it is more effective than directly supplying the source code of all methods within the software module to LLMs, as the latter could introduce excessive noise irrelevant to the target branch and confuse LLMs during test generation.
@@
Due to the non-negligible cost of utilizing LLMs, we define the usage scenario of TELPA as its activation when the existing (lightweight) test generation tool (e.g., Pynguin (Lukasczyk and Fraser, 2022)) fails to increase test coverage within a certain timeframe. In other words, TELPA is only employed for hard-to-cover branches to ensure cost-effectiveness. To enhance efficiency further, TELPA samples a diverse set of already-generated tests as counter-examples. By incorporating these counter-examples into the prompt, TELPA can instruct LLMs to generate tests that diverge from them, as these counter-examples have been recognized as ineffective for these hard-to-cover branches. In particular, the overall test generation process with TELPA is structured as a feedback-based process. This design allows for updating counter-examples based on the most recent coverage results, thereby enhancing test generation effectiveness.
@@
We conducted an extensive evaluation on 27 open-source Python projects, which have been widely used in previous studies (Lemieux et al., 2023; Lukasczyk and Fraser, 2022). Our results show that TELPA significantly outperforms both the state-of-the-art SBST tool (i.e., Pynguin (Lukasczyk and Fraser, 2022)) and the state-of-the-art LLM-based/LLM-enhanced techniques (i.e., CODAMOSA (Lemieux et al., 2023) and CHATTESTER (Yuan et al., 2024)). On average, TELPA achieves 34.10%, 25.93%, and 21.10% higher branch coverage than them respectively, given the same testing time budget. Our ablation study also confirms the contribution of each main component in TELPA, including object construction analysis for relieving the challenge arising from complex object construction, branch dependency analysis for relieving the challenge arising from intricate inter-procedural dependencies, counter-examples sampling and coverage-based feedback for improving the overall testing efficiency. Our investigation into various configurations of TELPA confirms that our designed usage scenario for TELPA strikes a balance between effectiveness and efficiency and TELPA consistently achieves higher coverage than other settings.
@@
We introduce TELPA, a novel LLM-based test generation technique designed to enhance the coverage of hard-to-cover branches through program-analysis-enhanced prompting.
TELPA can address the challenges arising from complex object construction and intricate inter-procedural dependencies well by incorporating both object construction and branch dependency analysis within TELPA.
@@
We present an extensive evaluation of TELPA across 27 open-source projects, demonstrating its effectiveness in comparison to the state-of-the-art SBST tool and two state-of-the-art LLM-based/LLM-enhanced techniques.
2. Motivating Example
In this section, we demonstrate the limitations of state-of-the-art automated test generation tools for Python, namely Pynguin (Lukasczyk and Fraser, 2022), CODAMOSA (Lemieux et al., 2023), and CHATTESTER (Yuan et al., 2024) with two real examples to motivate our study.
Figure 1 shows the simplified code snippet of the target method set_definition from the typesystem project111https://github.com/encode/typesystem. With a time budget of 20 minutes, Pynguin, CODAMOSA, and CHATTESTER all fail to reach line 6, namely the condition for the if statement in line 5 never returns true. In fact, to make the condition true, the generated test needs to first construct an Array object (line 3), and then assign a proper value to its items attribute. More specifically, items should be either a tuple or a list of Field objects (line 4), according to the type hints from the constructor of Array (line 13). Constructing such a sophisticated Array object is certainly challenging for automatic test generation techniques, Pynguin, CODAMOSA, and CHATTESTER are no exception.
Another type of hard-to-cover branches are caused by complicated inter-procedural dependencies, as illustrated in Figure 2. In this case, our target method is is_public_family from the apimd project222https://github.com/KmolYuan/apimd. Pynguin, CODAMOSA, and CHATTESTER are unable to guide the condition in line 3 to return true. The challenge here is that the outcome of the if condition depends on the method is_magic, which is defined elsewhere in the project and might further invoke other methods. Those complicated inter-procedural dependencies are often hard to understand and analyze for test generation techniques.
According to our analysis, our evaluation benchmark, which consists of 486 modules from 27 Python projects, contains 6,559 branches, of which 3,979 (60.7%) are classified as hard-to-cover. Similar challenges are also prevalent in Java, as evidenced by the Defects4J benchmark, which contains 9,391 branches, of which 5,494 (58.5%) are hard-to-cover. This highlights the prevalence of these challenges.
@@ 2. Motivating Example
First, during the test generation process, existing techniques (Lukasczyk and Fraser, 2022; Lemieux et al., 2023) normally try to construct minimal or over-simplified objects, without taking into account the real usage scenarios involving specific attributes. Second, for the inter-procedural dependencies, state-of-the-art techniques (Lukasczyk and Fraser, 2022; Lemieux et al., 2023) are not able to recursively analyze how dependent methods interact with variables from the target method, thus failing to interpret the concrete branch constraints. Third, existing LLM-based techniques (Lemieux et al., 2023; Yuan et al., 2024) typically provide either unnecessary or insufficient contexts for LLMs to generate tests. CODAMOSA provides the whole module as the context, where many parts of code are irrelevant to the target method and thus introduce noise to LLMs. CHATTESTER provides the class declaration, constructor signatures, relevant fields, and the target method as the context, which misses the broader context dependent to the target method.
@@
Based on these potential reasons behind the unsatisfactory performance of existing techniques, we propose TELPA, an LLM-based test generation technique leveraging program-analysis-enhanced prompting. TELPA will 1) conduct object construction analysis to learn real usage scenarios of objects and overcome the difficulty of constructing complex objects, 2) conduct branch dependency analysis to understand the behavior of complicated inter-procedural dependencies, and 3) supply relevant information to LLMs for test generation, avoiding imprecise context.
3. Approach
@@
The overall workflow of TELPA is illustrated in Figure 3. First, TELPA leverages existing test generation tools to generate a set of tests, which can reach those easy-to-cover branches. For ease of presentation, we call this step preceding test generation. Then, for the methods containing the branches that are still not covered, TELPA performs two types of program analyses: 1) object construction analysis: This analysis traces backward from the target method to identify sequences of method calls that lead to the construction of objects used as parameters. By capturing real usage scenarios, this analysis helps the LLM generate tests that construct valid objects with the necessary attributes; 2) branch dependency analysis: This analysis explores forward from the branch conditions of the target method to identify all methods that influence the outcome of the conditions. By providing precise contextual information about these dependencies, this analysis enables the LLM to generate tests that effectively cover hard-to-reach branches. To improve the testing efficiency, TELPA samples existing generated tests as counter-examples to guide LLMs to generate different tests, as they are shown ineffective to cover specific branches. These counter-examples, together with the program analysis results, are incorporated into the prompt for LLMs to generate new tests. The new tests will be executed and added into existing tests for future iterations. In the following, we will introduce TELPA in detail by using the example shown in Figure 4 for facilitating illustration.
3.1. Preceding Test Generation with Existing Tools
@@
As our goal is to reach those hard-to-cover branches, TELPA uses existing test generation tools for the initial testing task (for those easy-to-cover branches). This helps improve the cost-effectiveness of the overall testing process due to the cost of utilizing LLMs. Specifically, for a project under test, TELPA runs an existing test generation tool until it is unable to cover new branches within a pre-defined timeframe. Following previous work (Lemieux et al., 2023), we conducted a pilot study on a small benchmark and then set the timeframe to 2 minutes based on the observed results in TELPA. When executing these generated tests, their corresponding coverage information is also recorded. In this section, we mainly describe the idea behind TELPA, and the implementation details will be presented in Section 4.1.1.
3.2. Method-Invocation Analyses
@@
TELPA first collects the methods containing branches not covered by the preceding testing process, and then performs method invocation analyses on these methods to gather information needed for handling complex object construction and inter-procedural dependencies. This process is done in ascending order of coverage achieved for the methods, as we assume that methods with the lower coverage are more difficult to tackle and there is larger room for coverage improvement. Unlike existing LLM-based techniques (e.g., CODAMOSA (Lemieux et al., 2023)) that try to incorporate as much source code from the module as possible into prompts, we will only feed LLMs with the methods retrieved from method-invocation analyses as the context in order to reduce noise introduced by irrelevant code and alleviate computational burden of LLMs.
3.2.1. Object Construction Analysis